# Localizing Memorization in SSL Vision Encoders

### v1

@@ 8 Conv4-2 256 590080
Relying on insights from [Liu et al., 2021a], we calculate Equ. Equation 5 for the activations within UnitMem over ten augmentations since this has shown to yield a strong signal on the augmented data point. For convolutional feature maps, the activation of the unit is calculated as the average of all elements in the feature map. In ViTs, where we measure activation over fully-connected layers, we compute the activation per neuron and average across all patches of a given input. For example, for ViT Tiny encoder pretrained on CIFAR10, the input image of resolution 32x32 is patchified into 64 patches, each of size 4x4. Then, each patch is represented by a 192 dimensional embedding. The classification (CLS) embedding is prepended to the remaining 64 embeddings. Overall, we obtain 65 patches. The last fully connected layer has 192 neurons. For each neuron, we average its activations across the 65 patches. In the case of ViT Base, we have 768-dimensional embeddings and 197 patches for the input image of resolution 224x224.
@@ C.1 Additional Insights into UnitMem
UnitMem increases over training. First, we assess how UnitMem evolves over training of the SSL encoder. Therefore, we train a ResNet9 encoder using SimCLR on the CIFAR10 dataset for 800 epochs, using 120 warm-up epochs. Every five epochs, we measure the UnitMem. Our results in Fig. Figure 4 depict the average UnitMem of the ResNet9’s last convolutional layer.
@@
To assess the impact on using augmentation to implement Equ. Equation 5 for the calculation of our UnitMem has an impact on the reported results, we train two ResNet9 models on the CIFAR100 dataset, one using SimCLR, the other one using standard SL with cross entropy loss.
During training we rely on the standard augmentations for SL and SSL reported above. To measure memorization, we once use ten augmentations from the training augmentation set, and no augmentations otherwise and report the results in Fig. Figure 5. We find that while the trend of the reported memorization is equal in both settings, the UnitMem measured without augmentations remains constantly lower than when measured with augmentations. This suggests that when measuring UnitMem, it is important to use augmentations to avoid under-reporting of the memorization.
@@
Using the same ResNet9, trained with SimCLR on CIFAR100, we assess whether the number of data points that we use to calculate UnitMem (the size of $\mathcal{D^{\prime}}$) has an impact on the reported memorization. Then, we measure UnitMem using 100 random data point chosen one from each class in CIFAR100, 100 purely randomly chosen data points, and randomly chosen CIFAR100 data points. We present our findings in Fig. Figure 6. Our results highlight that all the lines are within each other’s standard deviation, indicating that there is no significant difference in the reported UnitMem, dependent on the make up of the dataset $\mathcal{D^{\prime}}$.
@@
We train a ResNet9 on CIFAR10 using SimCLR and measure UnitMem for the 300 most and 300 least memorized data points identified using SSLMem by [Wang et al., 2024]. The measurement of the two sets (most vs least memorized data points) is performed independently. Our results in Fig. Figure 7 show that the UnitMem calculated on the most memorized data points is significantly higher than on the least memorized data points (we verify the significance with a statistical $t$-test in Table 10.
@@
Our UnitMem metric is inspired by the class selectivity defined for SL by [Morcos et al., 2018b] which quantifies a unit’s discriminability between different classes, see Section D.1. Yet, we calculate the $\mu_{-max,u}$ in Equ. Equation 7 using the median on the other individual training data points’ activations while ClassSelectivity computes their equivalent of $\mu_{-max,u}$ using the mean on all other test classes’ activations.
@@
We show in Fig. Figure 8 over the 300 most and least memorized data points for a ResNet9 trained with SimCLR on CIFAR10 that using the median for UnitMem yields very similar results to using the mean.
@@
We additionally set out to study the impact of the augmentation set used to calculate UnitMem. Therefore, we calculate UnitMem on the ResNet9 trained on CIFAR10 using SimCLR using different augmentations sets. For SSL, we measure once with the standard training augmentations ("Normal"), with an independent set of augmentations of similar strength ("Independent"), with a weaker augmentation set for which we rely on the augmentations used to train the SL model ("Weak"), and an independent very strong set of augmentations modeled after MAE training and using a masking of 75% of the input image ("Masking"). Our results in Fig. Figure 9 depict the UnitMem over the last convolutional layer of the ResNet9 encoders.
@@
We compare the UnitMem score between corresponding layers of a ResNet50 pre-trained on ImageNet with SimCLR  and DINO, as well as for ViT-Base encoders pre-trained on ImageNet with DINO and MAE. We ensure that the number of epochs, batch sizes, training dataset sizes, and the resulting linear probing accuracies of the encoders are similar for direct comparability. Our results in Fig. Figure 11 depict the UnitMem of the last convolutional layer of the ResNet50, and the final block’s fully-connected layer in the ViT. The plot indicates that the different SSL frameworks applied to the same architecture with the same dataset yield similar memorization pattern.
@@ C.2 UnitMem Measures Memorization of Individual Data Points
To highlight that UnitMem reports memorization of individual data points rather than the a unit’s ability to recognize class-wide concepts, we designed an additional experiment. For the experiment, we rely on the class concept of "wheel" as an example. In the STL10 dataset, three classes have a concept wheel, namely Truck, Plane, Car. If UnitMem was to report simply a unit’s sensitivity to concepts of different classes (rather than individual data points), we would see a drop in UnitMem as we increase the percentage of data points with the concept wheel in the batch used to compute the metric. This is because then all data points should equally activate the unit, resulting in low memorization according to Equ. Equation 8.
@@ C.3 Additional Insights into LayerMem
C.5 Layer-based Memorization Across Different SSL Frameworks and Datasets
@@
C.6 Verification of Layer-Based Memorization
@@
C.7 Impact of Layer Replacement on Layer Memorization
@@
C.8 Layer Replacement for Single, Two, and Three Layers at a Time
@@ Appendix G Additional Results
$\ell_{1}$ $\ell_{2}$ Cos. Sim. Ang. Dist.
Layer LM $\Delta$LM LM $\Delta$LM LM $\Delta$LM LM $\Delta$LM
1 0.099 - 0.091 - 0.104 - 0.096 -
2 0.128 0.029 0.123 0.032 0.134 0.030 0.128 0.032
3 0.159 0.031 0.154 0.031 0.163 0.029 0.160 0.032
4 0.187 0.028 0.183 0.029 0.190 0.027 0.191 0.031
Res2222 0.192 0.005 0.185 0.002 0.193 0.003 0.193 0.002
5 0.221 0.029 0.212 0.027 0.220 0.027 0.222 0.029
6 0.256 0.035 0.246 0.034 0.256 0.036 0.259 0.037
7 0.289 0.033 0.276 0.030 0.288 0.032 0.293 0.034
8 0.325 0.036 0.308 0.032 0.321 0.033 0.328 0.035
Res6666 0.329 0.004 0.311 0.003 0.323 0.002 0.329 0.001
weak normal strong
Layer LM $\Delta$LM LM $\Delta$LM LM $\Delta$LM
1 0.092 - 0.091 - 0.089 -
2 0.123 0.031 0.123 0.032 0.120 0.031
3 0.154 0.031 0.154 0.031 0.150 0.030
4 0.184 0.030 0.183 0.029 0.178 0.028
Res2222 0.187 0.003 0.185 0.002 0.181 0.003
5 0.215 0.028 0.212 0.027 0.208 0.027
6 0.249 0.034 0.246 0.034 0.241 0.033
7 0.280 0.031 0.276 0.030 0.269 0.028
8 0.313 0.033 0.308 0.032 0.300 0.031
Res6666 0.315 0.002 0.311 0.003 0.302 0.002
Table 31: LayerMem (LM) and $\Delta$-LM under different augmentation sets used during training. We use the augmentations defined in lines 595ff. during training and metric calculation. The results show that larger augmentations lead to higher memorization. (ResNet9, CIFAR10, SimCLR).
Figure 12: UnitMem under different regularization strength used during training. We observe that with stronger regularization, memorization decreases, yet linear probing acc (LP. Acc.) decreases. (ResNet9, CIFAR10, SimCLR).
Figure 13: The most memorized samples per layer according to LayerMem. We observe consistency within adjacent layers that increases the further the layers are. (ResNet9, CIFAR10, SimCLR).
Layer 1 2 3 4 5 6 7 8
Overlap % 73 69 75 89 85 88 86 90
Layer 1 2 3 4 5 6 7 8
Cosine Similarity 0.797 0.844 0.823 0.811 0.779 0.805 0.847 0.881
Table 34: Cosine similarities between BatchNorm layer outputs for ResNet9 trained on CIFAR10 and STL10. We normalize the training data according to the ImageNet parameters and train the encoders using SimCLR. We calculate the cosine similarity over the weights ($\gamma$) and the bias ($\beta$) of the respective encoders’ trained BatchNorm layers.

### v2

@@ 4.1 Experimental Results and Observations
Variability and Consistency of Memorization cross Different Layers.
We use LayerMem to analyze the variability and consistency between the samples memorized by different layers in a ResNet9 vision encoder trained with CIFAR10 dataset. The results are shown in Table 3 and Figure 13 in section C.5. The overlap within the 100 most memorized samples between adjacent layers is usually high but decreases the further the layers are separated. Our statistical analysis to compare the similarity of the orderings within different layers’ most memorized samples using the Kendall’s rank correlation coefficient shows that while for closer layers, we manage to reject the null hypothesis (“no correlation”) with high statistical confidence (low p-value) which is not the case for further away layers.
@@ 6 Conclusions
Acknowledgments
The project on which this paper is based was funded by the Initiative and Networking Fund of the Helmholtz Association in the framework of the Helmholtz AI project call under the name „PAFMIM“, funding number ZT-I-PF-5-227. Responsibility for the content of this publication lies with the authors.
@@ 8 Conv4-2 256 590080
Relying on insights from [Liu et al., 2021a], we calculate  Equation 5 for the activations within UnitMem over ten augmentations since this has shown to yield a strong signal on the augmented data point. For convolutional feature maps, the activation of the unit is calculated as the average of all elements in the feature map. In ViTs, where we measure activation over fully-connected layers, we compute the activation per neuron and average across all patches of a given input. For example, for ViT Tiny encoder pretrained on CIFAR10, the input image of resolution 32x32 is patchified into 64 patches, each of size 4x4. Then, each patch is represented by a 192 dimensional embedding. The classification (CLS) embedding is prepended to the remaining 64 embeddings. Overall, we obtain 65 patches. The last fully connected layer has 192 neurons. For each neuron, we average its activations across the 65 patches. In the case of ViT Base, we have 768-dimensional embeddings and 197 patches for the input image of resolution 224x224.
@@ C.1 Additional Insights into UnitMem
UnitMem increases over training. First, we assess how UnitMem evolves over training of the SSL encoder. Therefore, we train a ResNet9 encoder using SimCLR on the CIFAR10 dataset for 800 epochs, using 120 warm-up epochs. Every five epochs, we measure the UnitMem. Our results in Figure 4 depict the average UnitMem of the ResNet9’s last convolutional layer.
@@
To assess the impact on using augmentation to implement Equation 5 for the calculation of our UnitMem has an impact on the reported results, we train two ResNet9 models on the CIFAR100 dataset, one using SimCLR, the other one using standard SL with cross entropy loss.
During training we rely on the standard augmentations for SL and SSL reported above. To measure memorization, we once use ten augmentations from the training augmentation set, and no augmentations otherwise and report the results in Figure 5. We find that while the trend of the reported memorization is equal in both settings, the UnitMem measured without augmentations remains constantly lower than when measured with augmentations. This suggests that when measuring UnitMem, it is important to use augmentations to avoid under-reporting of the memorization.
@@
Using the same ResNet9, trained with SimCLR on CIFAR100, we assess whether the number of data points that we use to calculate UnitMem (the size of $\mathcal{D^{\prime}}$) has an impact on the reported memorization. Then, we measure UnitMem using 100 random data point chosen one from each class in CIFAR100, 100 purely randomly chosen data points, and randomly chosen CIFAR100 data points. We present our findings in Figure 6. Our results highlight that all the lines are within each other’s standard deviation, indicating that there is no significant difference in the reported UnitMem, dependent on the make up of the dataset $\mathcal{D^{\prime}}$.
@@
We train a ResNet9 on CIFAR10 using SimCLR and measure UnitMem for the 300 most and 300 least memorized data points identified using SSLMem by [Wang et al., 2024]. The measurement of the two sets (most vs least memorized data points) is performed independently. Our results in Figure 7 show that the UnitMem calculated on the most memorized data points is significantly higher than on the least memorized data points (we verify the significance with a statistical $t$-test in Table 11.
@@
Our UnitMem metric is inspired by the class selectivity defined for SL by [Morcos et al., 2018b] which quantifies a unit’s discriminability between different classes, see Section D.1. Yet, we calculate the $\mu_{-max,u}$ in  Equation 7 using the median on the other individual training data points’ activations while ClassSelectivity computes their equivalent of $\mu_{-max,u}$ using the mean on all other test classes’ activations.
@@
We show in Figure 8 over the 300 most and least memorized data points for a ResNet9 trained with SimCLR on CIFAR10 that using the median for UnitMem yields very similar results to using the mean.
@@
We additionally set out to study the impact of the augmentation set used to calculate UnitMem. Therefore, we calculate UnitMem on the ResNet9 trained on CIFAR10 using SimCLR using different augmentations sets. For SSL, we measure once with the standard training augmentations ("Normal"), with an independent set of augmentations of similar strength ("Independent"), with a weaker augmentation set for which we rely on the augmentations used to train the SL model ("Weak"), and an independent very strong set of augmentations modeled after MAE training and using a masking of 75% of the input image ("Masking"). Our results in Figure 9 depict the UnitMem over the last convolutional layer of the ResNet9 encoders.
@@
Figure 11: Different weight decay
Stronger weight decay reduces memorization.
To analyze how training weight decay affects the final encoder’s UnitMem, we train a ResNet9 using SimCLR on CIFAR10 with three different levels of weight decay. Our results in Figure 11 show that stronger weight decay yields lower memorization, yet also decreases linear probing accuracy.
@@
We compare the UnitMem score between corresponding layers of a ResNet50 pre-trained on ImageNet with SimCLR and DINO, as well as for ViT-Base encoders pre-trained on ImageNet with DINO and MAE. We ensure that the number of epochs, batch sizes, training dataset sizes, and the resulting linear probing accuracies of the encoders are similar for direct comparability. Our results in Figure 12 depict the UnitMem of the last convolutional layer of the ResNet50, and the final block’s fully-connected layer in the ViT. The plot indicates that the different SSL frameworks applied to the same architecture with the same dataset yield similar memorization pattern.
@@ C.2 UnitMem Measures Memorization of Individual Data Points
To highlight that UnitMem reports memorization of individual data points rather than the a unit’s ability to recognize class-wide concepts, we designed an additional experiment. For the experiment, we rely on the class concept of "wheel" as an example. In the STL10 dataset, three classes have a concept wheel, namely Truck, Plane, Car. If UnitMem was to report simply a unit’s sensitivity to concepts of different classes (rather than individual data points), we would see a drop in UnitMem as we increase the percentage of data points with the concept wheel in the batch used to compute the metric. This is because then all data points should equally activate the unit, resulting in low memorization according to Equation 8.
@@ C.3 Additional Insights into LayerMem
C.5 Visualization for Variability and Consistency of Memorization cross Different Layers.
We present the top 10 most memorized samples of each layer for the ResNet9 vision encoder trained with the CIFAR10 dataset in Figure 13. The results show that the overlap within the top 10 most memorized samples between adjacent layers is usually high but decreases the further the layers are separated. This aligns with the results of overlap rate and Kendall’s Tau test reported in Table 3.
Figure 13: The most memorized samples per layer according to LayerMem.
C.6 Layer-based Memorization Across Different SSL Frameworks and Datasets
@@
C.7 Verification of Layer-Based Memorization
@@
Statistics of Batch-Norm layer for different datasets.
Batch-norm layers between different datasets might have different statistics. This could impact the downstream performance. To investigate the changes, we measured the cosine similarity between the weights and biases of the batch-norm layers for two encoders (trained on CIFAR10 and STL10, respectively). The results in Table 25 show a high per-layer cosine similarity (average over all layers=0.823). This suggests that the statistics are similar, hence, no adjustment is required. We hypothesize that the similarity stems from the fact that the data distributions are similar and that we normalize both input datasets according to the ImageNet normalization parameters.
Table 25: Cosine similarities between batch-norm layer outputs for ResNet9 trained on CIFAR10 and STL10. We normalize the training data according to the ImageNet parameters and train the encoders using SimCLR. We calculate the cosine similarity over the weights ($\gamma$) and the bias ($\beta$) of the respective encoders’ trained batch-norm layers.
Layer 1 2 3 4 5 6 7 8
Cosine Similarity 0.797 0.844 0.823 0.811 0.779 0.805 0.847 0.881
C.8 LayerMem with Different Distance Metrics
In addition to the $\ell_{2}$ distance, we also used 3 other distance metrics ($\ell_{1}$, cosine similarity, and angular distance) to evaluate the stability of LayerMem. Our results in Table 26 highlight that 1) the memorization scores are very similar, independent of the choice of the distance metric, and 2) the most memorizing layers according to and $\Delta$LayerMem are the same over all metrics. This suggests that our findings are independent of the choice of distance metric.
$\ell_{1}$ $\ell_{2}$ Cos. Sim. Ang. Dist.
Layer LM $\Delta$LM LM $\Delta$LM LM $\Delta$LM LM $\Delta$LM
1 0.099 - 0.091 - 0.104 - 0.096 -
2 0.128 0.029 0.123 0.032 0.134 0.030 0.128 0.032
3 0.159 0.031 0.154 0.031 0.163 0.029 0.160 0.032
4 0.187 0.028 0.183 0.029 0.190 0.027 0.191 0.031
Res2222 0.192 0.005 0.185 0.002 0.193 0.003 0.193 0.002
5 0.221 0.029 0.212 0.027 0.220 0.027 0.222 0.029
6 0.256 0.035 0.246 0.034 0.256 0.036 0.259 0.037
7 0.289 0.033 0.276 0.030 0.288 0.032 0.293 0.034
8 0.325 0.036 0.308 0.032 0.321 0.033 0.328 0.035
Res6666 0.329 0.004 0.311 0.003 0.323 0.002 0.329 0.001
C.9 LayerMem with Different Augmentation Strength
The results, reported in Table 27 highlight that stronger training augmentations reduce LayerMem.
Table 27: LayerMem (LM) and $\Delta$-LM under different augmentation sets used during training. We use the augmentations defined in Appendix B during training and metric calculation. The results show that stronger augmentations reduce memorization. (ResNet9, CIFAR10, SimCLR).
weak normal strong
Layer LM $\Delta$LM LM $\Delta$LM LM $\Delta$LM
1 0.092 - 0.091 - 0.089 -
2 0.123 0.031 0.123 0.032 0.120 0.031
3 0.154 0.031 0.154 0.031 0.150 0.030
4 0.184 0.030 0.183 0.029 0.178 0.028
Res2222 0.187 0.003 0.185 0.002 0.181 0.003
5 0.215 0.028 0.212 0.027 0.208 0.027
6 0.249 0.034 0.246 0.034 0.241 0.033
7 0.280 0.031 0.276 0.030 0.269 0.028
8 0.313 0.033 0.308 0.032 0.300 0.031
Res6666 0.315 0.002 0.311 0.003 0.302 0.002
C.10 LayerMem with Different Initialization of Trainable parameters
We performed an additional experiment where we trained encoders f and g independently with a different random seed (yielding f’ and g’) to study how random initialization of trainable parameters can affect the memorization of final vision encoder.The results are reported in Table 28. We compared the overlap in most memorized samples between encoder f (from the paper) and f’. The results (Table 4, attached PDF) show that overlap is overall high (min. 69% in layer 2) and increases in the later layers (max. 90%, final layer).
Layer 1 2 3 4 5 6 7 8
Overlap % 73 69 75 89 85 88 86 90
C.11 Impact of Layer Replacement on Layer Memorization
@@
C.12 Layer Replacement for Single, Two, and Three Layers at a Time
@@ Appendix G Additional Results
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We introduce our main contributions and key findings from line 5 to line 19 in the abstract and line 73 to line 80 in Section 1
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss our Limitations in Appendix F
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: All related results are clearly stated and referenced in either the main paper or the appendix.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The detailed experimental setup is introduced in Appendix B and related source code is uploaded to open-review.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: All related source code is uploaded to open-review and experiments are conducted on open source datasets.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The detailed experimental setup is introduced in Appendix B.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: For the average results of all multiple experiments in the paper, we report the standard deviation in the tables and draw the error bar used to represent the standard deviation in the figures.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The hardware usage is introduced in Appendix B
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  45. 10.
    Broader Impacts
  46. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  47. Answer: [N/A]
  48. Justification: In this work, we introduce two metrics for locating memorization in SSL vision encoders, as well as some key findings that result from experimenting with our metrics. No potential direct social impact is expected.
  49. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  50. 11.
    Safeguards
  51. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  52. Answer: [N/A]
  53. Justification: All datasets used to train models for this paper are safe public datasets, including ImageNet ILSVRC-2012 [Russakovsky et al., 2015], CIFAR10 [Krizhevsky et al., 2009], CIFAR100 [Krizhevsky et al., 2009], SVHN [Netzer et al., 2011], and STL10 [Coates et al., 2011] (This is also introduced in Appendix B).
  54. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  55. 12.
    Licenses for existing assets
  56. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  57. Answer: [Yes]
  58. Justification: All models and datasets usage is detailed introduced in Appendix B. We totally abbey the terms of use of these public sources. All other code works are done by ourselves with no potential risks of licenses.
  59. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  60. 13.
    New Assets
  61. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  62. Answer: [Yes]
  63. Justification: We provide all our code for metrics mentioned in our paper to open review and write a detailed Readme file for the usage of our code.
  64. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  65. 14.
    Crowdsourcing and Research with Human Subjects
  66. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  67. Answer: [N/A]
  68. Justification:
  69. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  70. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  71. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  72. Answer: [N/A]
  73. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.